OpenStudy (anonymous):

(sqrt 7 - sqrt 3)( sqrt 7 + sqrt 3)

OpenStudy (anonymous):

7-3... 4?

OpenStudy (anonymous):

7 -3=4 because (a-b)(a+b)= a^2 - b^2